# Dynamic Multi-Hop HARQ (DMH-HARQ): Reliable and Open Latency-Constrained Wireless Transport Network

### v3

@@ IV-B Multi-Hop Link Reliability with Perfect Global CSI
Based on the error model above, we start our analysis to the optimization problem (III-C) with a simplified case where perfect CSI of all hops $i\in\mathcal{I}$ is available at every node. For the convenience of notation, in the analyses below we let
$\xi_{\mathrm{max}}(s)\triangleq\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S},$ (10)
@@
$n_{\mathrm{opt}}(s)\triangleq\arg\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S}.$ (11)
@@
$\begin{split}&\xi_{\mathrm{max}}(s)\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))\right\}\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\max\limits_{n_{S(L(\mathfrak{T}_{s}))}}\xi(L(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\max\limits_{n_{S(R(\mathfrak{T}_{s}))}}\xi(R(\mathfrak{T}_{s}))\right\}\\ =&\dots\end{split}$ (12)
which forms a Bellman equation. Under the error model (8), we have the following conclusions:
@@
IV-C Multi-Hop Link Reliability with Imperfect CSI
$\overline{\varepsilon}_{s^{\prime}}=\int\limits_{0}^{+\infty}\varepsilon_{s^{\prime}}f_{G}(g_{i_{s}^{\prime}})\text{d}g_{i_{s^{\prime}}}.$ (13)
V Algorithm Design
V-A Optimal DMH-HARQ: Integer Dynamic Programming
From Lemma 1 we know that when applying Type I HARQ without combining, the DMH-HARQ scheduling is a typical Markov decision process (MDP), where the expected utility $\xi_{s}$ from any certain status $s$ is independent from its historical states or actions, but only relying on the remaining time $t_{s}$ and the blocklength allocation policy $n_{s}$. Such kind of problems can be generally solved by recursive approaches that gradually evaluate and improve the decision policy by traversing the entire feasible region of states, typical examples are policy iteration and value iteration. However, such approaches are challenged by two issues when dealing with Eq. (12). First, (8) provides no closed-form analytical solution $n_{s}$ for an arbitrary $\varepsilon_{s}$. Second, with $n_{s}\in\mathbb{R}^{+}$, both the feasible region $\mathcal{S}$ and the set of possible actions becomes infinite, making it impossible to traverse the solution space within limited time.
{strip}
@@ V-A Optimal DMH-HARQ: Integer Dynamic Programming
             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}_{\mathrm{local}}.\mathrm{find}\left([t_{s},i_{s}]\right)$
@@
             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
@@
17             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+[1-\varepsilon_{s}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})\}$
@@
19             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\varepsilon_{s}(n_{\mathrm{opt}})]\textnormal{{BestSucc!}}(t_{s}-n_{\mathrm{opt}}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})$
@@
30             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}.\mathrm{find}\left([t_{s},i_{s}]\right)$
@@
32             $[\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
@@
39             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})\}$
@@
41             ${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})$
@@ 51             $\mathfrak{T}_{s}\leftarrow\emptyset$
55             ${}~{}~{}~{}~{}~{}~{}~{}\left.\textnormal{{ScheduleTreeGen}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,0,0,\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}})\right)$
@@
Nevertheless, it shall be remarked that the real-value relaxation $n_{s}\in\mathbb{R}^{+}$ was taken, like in classical FBL works, only for the convenience of analysis. Indeed, representing the blocklength per transmission in the unit of channel use, the final solution $n_{\mathrm{opt}}(s)$ can only take integer values from $\mathbb{N}^{+}$. Meanwhile, to apply FBL approaches while guaranteeing to meet the stringent latency requirement, as referred earlier, $n_{s}$ is strictly constrained by (III-C). Thus, both the action set and the feasible region of status are usually of reasonable sizes, enabling us to apply classical integer DP techniques to directly solve the global optimum (the feasibility and optimality of this approach are discussed in [33]). A typical implementation is to recursively solve $n_{\mathrm{opt}}(s)$ from the last hop where $i_{s}=1$ on, hop-by-hop onto the first one where $i_{s}=I$. The recursive algorithm is essentially accompanied with LUT to store the optimal achievable utility over the solution space, as described by Algorithm 1:
@@ V-B Hop-by-Hop CSI Update and Reschedule
Note that by running Algorithm 1 at node $i\in\mathcal{I}$, the optimal schedule $\mathfrak{T}_{s}$ is solved based on the perfect CSI of the $i^{\mathrm{th}}$ hop and statistical CSI of all hops thereafter, which shall not be adopted by any other node. Indeed, upon a successful decoding of the message received from its predecessor node, each node $i\in\mathcal{I}$ shall individually solve its own optimal schedule $\mathfrak{T}_{s}$ regarding the instantaneous system state $s$ and its real-time channel measurement $g_{i}$ of the $i^{\mathrm{th}}$ hop. The complete procedure can be briefly summarized as in Algorithm 2.
@@
2Initialization: $[g_{1},g_{2},\dots g_{I}],f_{G},\varepsilon_{\mathrm{max}},T_{\mathrm{max}},T_{\mathrm{b}},T_{\mathrm{l}}$
@@ 9             Update $s$ regarding the transmission result
Algorithm 2 $I$-hop DMH-HARQ with CSI update
@@ V-C Computational Complexity Analysis
As described above, Algorithm 1 implements a recursive DP approach to compute the integer dynamic program. The recursive algorithm considers the DMH-HARQ as an $I$-stage DP problem, each stage representing a hop. Meanwhile, each hop-level sub-problem that starting with state $s$ makes also an $\mathcal{O}(n_{\mathrm{max},s})$-stage DP problem itself, each stage representing an available channel use in the schedule.
According to the computational complexity analysis of the CLARQ algorithm [33], which has the same structure as each hop-level sub-problem, we know that the time complexity of each hop-level sub-problem is $\mathcal{O}(n_{\mathrm{max},s})$. Since for each $s\in\mathcal{S}$ hop, $n_{\text{max},s}$ is upper-bounded by $n_{\mathrm{MAX}}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(I-1)}}{Mf_{\mathrm{s}}}$, the time complexity of solving the optimal DMH-HARQ at the $i^{\mathrm{th}}$ last node is ${\mathcal{O}\left(n_{\mathrm{MAX}}^{i}\right)}$. Taking the hop-by-hop execution of Algorithm 2 into account, the overall time complexity of the DMH-HARQ approach over the $I$-hop relay chain is $\mathcal{O}\left[\sum\limits_{i=1}^{I}n_{\mathrm{MAX}}^{i}\right]=\mathcal{O}\left(n_{\mathrm{MAX}}^{I}\right)$.
In realistic scenarios of deployment, such a time complexity can critically challenge the online computation of optimal DMH-HARQ schedule, especially due to the time-varying channel gains $g_{i}$, which leads to high delay and significant power consumption. To address this issue, it is a practical solution to implement a LUT in each node, which that contains a set of offline solved dictionaries regarding various channel conditions. Since the operation of searching for a specific entry in a dictionary-based LUT has only a time complexity of $\mathcal{O}(1)$, the devices can be rapidly adapted to the appropriate specification w.r.t. the real-time channel measurement, leading to an overall time complexity as low as $\mathcal{O}(I)$ for the entire $I$-hop DMH-HARQ process.
@@ VI Numerical Evaluation
$M$ $2$ Modulation order, 2 for BPSK
@@
[…]
$\varepsilon_{\mathrm{max}}$ $0.5$ Maximal allowed PER per transmission
[…]
$I$ $4$ Default number of hops
$\mathbf{I}$$\{2,3,\dots,10\}$ Test range of hops number
$L_{\text{f},i}$ $\sim\mathrm{Rice}(0.5,1)$ Small-scale fading with a consistent LoS path
[…]
@@
$n_{\mathrm{MC}}$ $1~{}000~{}000$ Runs per Monte-Carlo test
@@ VI-A Baseline Methods
    Listening-based cooperative ARQ: which was proposed by Goel et al. in [36]. Similar to the naive static HARQ approach, every hop is assigned with the same length of sub-frame and pre-scheduled with the same number of (re)transmissions and coding rate. However, if the DF on some hop $i$ succeeds before using up all the pre-scheduled (re)transmission slots assigned to it, the DF on hop $i+1$ will immediately be triggered, inheriting all the remaining time resource from its predecessor hop. In our simulation, each hop is pre-scheduled with $2$ (re)transmission slots, applying Type-I HARQ without combining.
The principles of both static HARQ-IR and listening-based cooperative ARQ are illustrated in Fig. 2 for comparison with Fig. 1. Note that the naive and optimal schemes of static HARQ-IR only distinguish from each other by weather the sub-frames are of of same length over all hops.
@@ VI-B Simulation Setup
The setup of our simulation campaign is detailed in Tab. I. Among all the parameters, there are four that are of particular interest to us to test the sensitivity of our approach, namely the total time frame length $T_{\mathrm{max}}$, the number of hops $I$, the decoding/feedback delay $T_{\mathrm{l}}$, and the mean channel gain $\overline{g}$. Each of these key parameters is assigned with a default value and a test range. When benchmarking the approaches regarding one key parameter over its test range, all the rest key parameters are specified to their default values. For every individual specification, $n_{\mathrm{MC}}=1~{}000~{}000$ runs of Monte-Carlo test are conducted, each run with an independently generated set of channel conditions for the $I$ hops, with which all approaches are tested.

### v4

@@ IV-B Multi-Hop Link Reliability with Perfect Global CSI
Based on the error model above, we start our analysis to the optimization problem (6) with a simplified case where perfect CSI of all hops $i\in\mathcal{I}$ is available at every node. For the convenience of notation, in the analyses below we let
$\xi_{\mathrm{max}}(s)\triangleq\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S},$ (6l)
@@
$n_{\mathrm{opt}}(s)\triangleq\arg\max\limits_{n:\mathcal{S}\to\mathbb{R}^{+}}\xi(\mathfrak{T}_{s}),\qquad\forall s\in\mathcal{S}.$ (6m)
@@
$\begin{split}&\xi_{\mathrm{max}}(s)\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\xi_{\mathrm{max}}(s_{\mathrm{l}}(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\xi_{\mathrm{max}}(s_{\mathrm{r}}(\mathfrak{T}_{s}))\right\}\\ =&\max\limits_{n_{s}}\left\{\varepsilon_{s}\max\limits_{n_{S(L(\mathfrak{T}_{s}))}}\xi(L(\mathfrak{T}_{s}))+(1-\varepsilon_{s})\max\limits_{n_{S(R(\mathfrak{T}_{s}))}}\xi(R(\mathfrak{T}_{s}))\right\}\\ =$$\dots\end{split}$ (6n)
which forms a Bellman equation. Under the error model (6i), we have the following conclusions:
@@
     $\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}_{\mathrm{local}}.\mathrm{find}\left([t_{s},i_{s}]\right)$
@@
     $\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
@@
17     $1-\varepsilon_{s}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})\}$
@@
19     $1-\varepsilon_{s}(n_{\mathrm{opt}})]\textnormal{{BestSucc!}}(t_{s}-n_{\mathrm{opt}}T_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D}_{\mathrm{succ}})$
@@
30     $\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow\mathcal{D}.\mathrm{find}\left([t_{s},i_{s}]\right)$
@@
32     $\xi_{\mathrm{opt}},n_{\mathrm{opt}}]\leftarrow[0,0]$
@@
39     $1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})\}$
@@
41     $1-\overline{\varepsilon}(n)]\textnormal{{BestSucc!}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,f_{G},\varepsilon_{\mathrm{max}},\mathcal{D})$
@@ 51     $\mathfrak{T}_{s}\leftarrow\emptyset$
55     $\left.\textnormal{{ScheduleTreeGen}}(t_{s}-nT_{\mathrm{b}}-T_{\mathrm{l}},i_{s}-1,0,0,\mathcal{D}_{\mathrm{local}},T_{\mathrm{b}},T_{\mathrm{l}})\right)$
@@
IV-C Multi-Hop Link Reliability with Imperfect CSI
$\overline{\varepsilon}_{s^{\prime}}=\int\limits_{0}^{+\infty}\varepsilon_{s^{\prime}}f_{G}(g_{i_{s}^{\prime}})\text{d}g_{i_{s^{\prime}}}.$ (6o)
V Algorithm Design
V-A Optimal DMH-HARQ: Integer Dynamic Programming
From Lemma 1 we know that when applying Type I HARQ without combining, the DMH-HARQ scheduling is a typical Markov decision process (MDP), where the expected utility $\xi_{s}$ from any certain status $s$ is independent from its historical states or actions, but only relying on the remaining time $t_{s}$ and the blocklength allocation policy $n_{s}$. Such kind of problems can be generally solved by recursive approaches that gradually evaluate and improve the decision policy by traversing the entire feasible region of states, typical examples are policy iteration and value iteration. However, such approaches are challenged by two issues when dealing with Eq. (6n). First, (6i) provides no closed-form analytical solution $n_{s}$ for an arbitrary $\varepsilon_{s}$. Second, with $n_{s}\in\mathbb{R}^{+}$, both the feasible region $\mathcal{S}$ and the set of possible actions becomes infinite, making it impossible to traverse the solution space within limited time.
Nevertheless, it shall be remarked that the real-value relaxation $n_{s}\in\mathbb{R}^{+}$ was taken, like in classical FBL works, only for the convenience of analysis. Indeed, representing the blocklength per transmission in the unit of channel use, the final solution $n_{\mathrm{opt}}(s)$ can only take integer values from $\mathbb{N}^{+}$. Meanwhile, to apply FBL approaches while guaranteeing to meet the stringent latency requirement, as referred earlier, $n_{s}$ is strictly constrained by (6a). Thus, both the action set and the feasible region of status are usually of reasonable sizes, enabling us to apply classical integer DP techniques to directly solve the global optimum (the feasibility and optimality of this approach are discussed in [36]). A typical implementation is to recursively solve $n_{\mathrm{opt}}(s)$ from the last hop where $i_{s}=1$ on, hop-by-hop onto the first one where $i_{s}=I$. The recursive algorithm is essentially accompanied with LUT to store the optimal achievable utility over the solution space, as described by Algorithm 1:
@@ V-B Hop-by-Hop CSI Update and Reschedule
Note that by running Algorithm 1 at node $i\in\mathcal{I}$, the optimal schedule $\mathfrak{T}_{s}$ is solved based on the perfect CSI of the $i^{\mathrm{th}}$ hop and statistical CSI of all hops thereafter, which shall not be adopted by any other node. Indeed, upon a successful decoding of the message received from its predecessor node, each node $i\in\mathcal{I}$ shall individually solve its own optimal schedule $\mathfrak{T}_{s}$ regarding the instantaneous system state $s$ and its real-time channel measurement $g_{i}$ of the $i^{\mathrm{th}}$ hop. The complete procedure can be briefly summarized as in Algorithm 2. Note that the total number of hops $I$ is not endogenous to the algorithm, but only passed to the first node to initialize the state $s$. This allows our approach to be flexibly applied in the practical scenarios, where the number of hops is dynamically determined by the specific service function chain and the network topology.
@@
2Initialization: $g_{1},g_{2},\dots g_{I}],f_{G},\varepsilon_{\mathrm{max}},T_{\mathrm{max}},T_{\mathrm{b}},T_{\mathrm{l}}$
@@ 9     Update $s$ regarding the transmission result
Algorithm 2 II-hop DMH-HARQ with CSI update
@@ V-C Computational Complexity Analysis
As described above, Algorithm 1 implements a recursive DP approach to compute the integer dynamic program. The recursive algorithm considers the DMH-HARQ as an $I$-stage DP problem, each stage representing a hop. Meanwhile, each hop-level sub-problem that starting with state $s$ makes also an $\mathcal{O}(n_{\mathrm{max},s})$-stage DP problem itself, each stage representing an available channel use in the schedule. With any limited $T_{\max}$, maximum of level of stage is finite and upper-bounded, i.e., $I_{\max}\leqslant\frac{T_{\max}}{Mf_{s}n_{\min,i_{s}}}$, where $n_{\min,i_{s}}$ is the minimum blocklength defined in (6g). Therefore, the convergence of the algorithm is guaranteed.
According to the computational complexity analysis of the CLARQ algorithm [36], which has the same structure as each hop-level sub-problem, we know that the time complexity of each hop-level sub-problem is $\mathcal{O}(n_{\mathrm{max},s})$. Since for each $s\in\mathcal{S}$ hop, $n_{\text{max},s}$ is upper-bounded by $n_{\mathrm{MAX}}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(I-1)}}{Mf_{\mathrm{s}}}$, the time complexity of solving the optimal DMH-HARQ at the $i^{\mathrm{th}}$ last node is $\mathcal{O}\left(n_{\mathrm{MAX}}^{i}\right)$}. Taking the hop-by-hop execution of Algorithm 2 into account, the overall time complexity of the DMH-HARQ approach over the $I$-hop relay chain is $\mathcal{O}\left[\sum\limits_{i=1}^{I}n_{\mathrm{MAX}}^{i}\right]=\mathcal{O}\left(n_{\mathrm{MAX}}^{I}\right)$, which is significantly higher than to optimize the schedule of conventional static HARQ approaches, i.e. $\mathcal{O}\left[\log\left(n_{\mathrm{MAX}}\right)\right]$.
In realistic scenarios of deployment, such a time complexity can critically challenge the online computation of optimal DMH-HARQ schedule, especially due to the time-varying channel gains $g_{i}$, which leads to high delay and significant power consumption. To address this issue, it is a practical solution to implement a LUT in each node, which that contains a set of offline solved dictionaries regarding various channel conditions, i.e., the integer DP algorithm is pre-executed offline, only the LUTs must be deployed at different relaying nodes. Since the operation of searching for a specific entry in a dictionary-based LUT has only a time complexity of $\mathcal{O}(1)$, the devices can be rapidly adapted to the appropriate specification w.r.t. the real-time channel measurement, leading to an overall time complexity as low as $\mathcal{O}(I)$ for the entire $I$-hop DMH-HARQ process.
@@ VI Numerical Evaluation
$Q_{\mathrm{m}}$ 11 Modulation order, 1 for BPSK
$P$ $1\text{\,}\mathrm{mW}$ Transmission power
NN $1\text{\,}\mathrm{mW}$ Noise power
$d$ $16\text{\,}\mathrm{bit}$ Payload of each message
$\varepsilon_{\mathrm{max}}$ $0.5$ Maximal allowed PER per transmission
$T_{\mathrm{max}}$ $1\text{\,}\mathrm{ms}$ Default time frame length
II 44 Default number of hops
$\mathbf{I}$ $\{2,3,\dots,10\}$ Test range of hops number
$L_{\text{f},i}$ $\sim\mathrm{Rice}(0.5,1)$ Small-scale fading with a consistent line-of-sight (LoS) path
@@
$\overline{g}$ $0\text{\,}\mathrm{dB}$ Mean channel gain
@@
$n_{\mathrm{MC}}$ 1000000{1~000~000} Runs per Monte-Carlo test
@@ VI-A Baseline Methods
    Listening-based cooperative ARQ: which was proposed by Goel et al. in [12]. More specifically, here we consider its fully-cumulative scheme which outperforms other variants. Similar to the naive static HARQ approach, every hop is assigned with the same length of sub-frame and pre-scheduled with the same number of (re)transmissions and coding rate. However, if the DF on some hop $i$ succeeds before using up all the pre-scheduled (re)transmission slots assigned to it, the DF on hop $i+1$ will immediately be triggered, inheriting all the remaining time resource from its predecessor hop. In our simulation, each hop is pre-scheduled with $2$ (re)transmission slots, applying Type-I HARQ without combining.
The principles of both static HARQ-IR and listening-based cooperative ARQ are illustrated in Fig. 3 for comparison with Fig. 2. Note that the naive and optimal schemes of static HARQ-IR only distinguish from each other by whether the sub-frames are of of same length over all hops.
@@ VI-B Simulation Setup
The setup of our simulation campaign is detailed in Tab. I. It is worth remarking that we consider here the E2E delay $T_{\mathrm{max}}$ no longer than $1\text{\,}\mathrm{ms}$. Considering the low mobility of transceivers of wireless backhaul channels, this setup ensures that the delay over each single hop is lower than the coherence time of LoS channels in the centimeter wave range [13, 14], which is commonly used for such channels. Among all the parameters, there are four that are of particular interest to us to test the sensitivity of our approach, namely the total time frame length $T_{\mathrm{max}}$, the number of hops $I$, the decoding/feedback delay $T_{\mathrm{l}}$, and the mean channel gain $\overline{g}$. Each of these key parameters is assigned with a default value and a test range. When benchmarking the approaches regarding one key parameter over its test range, all the rest key parameters are specified to their default values. For every individual specification, $n_{\mathrm{MC}}={1~000~000}$ runs of Monte-Carlo test are conducted, each run with an independently generated set of channel conditions for the $I$ hops, with which all approaches are tested.